OpenStudy (anonymous):

Teresa has a rectangular poster that is 16 centimeters long and 10 centimeters wide. What is the area of the poster in square meters?

OpenStudy (anonymous):

area of a rectangle = length * breadth.So the area of the poster is 160 sq. cm

OpenStudy (anonymous):

the area of any rectangular shape can be found by multiplying the two sidelenghts by eachother. 16 long times 10 wide is 160 cm SQUARED. (all area is measured as cm^2)

OpenStudy (anonymous):

don't forget your conversion factor to get it into sq meters

OpenStudy (anonymous):

sorry I should have read the question completely.Since you want it in sq meters.It is probably easier if you first convert everything into meters. So 16 cm = 16/100 meters = 0.16 m 10 cm = 10/100 meters = 0.1 m so the area = 0.16 * 0.1 = .016 sq. meters

OpenStudy (anonymous):

oops. good call. ED: dont count this one as one of my ten